plz ans - calculate how much heat energy is required to raise the temperature of a cube of copper of side 2cm. from 300 to 320k. given: density=8933kg/m cube , atomic weight=63.55, cp=24j/k/mole

plz ans -


calculate how much heat energy is required to raise the temperature of a cube of copper of side 2cm. from 300 to 320k. given: density=8933kg/m cube , atomic weight=63.55, cp=24j/k/mole

Grade:12th Pass

1 Answers

Akash Kumar Dutta
11 years ago

Dear Roopa,

your question is based on a simple formula,
heat energy required = mcT
mass=d.v=8933 . (.02)^3 = .0712 kg = 71.2 gm
mo of moles =  71.2 / 63.55 = 1.12 moles
c=24 j/k-mole.
T=20
H= (1.12) . 24 . 20
  =537.6 Joules (ANS)

Regards.